OpenStudy (lgbasallote):

find the parabola with equation y = ax^2 + bx whose tangent line at (1,1) has equation y = 3x-2

OpenStudy (anonymous):

first we find b

OpenStudy (ash2326):

@lgbasallote I'd written so much but chrome showed me Aww snapp

OpenStudy (lgbasallote):

AWWWW...we need to report this :(

OpenStudy (anonymous):

3x - 2 = 2ax + b

OpenStudy (anonymous):

at (1,1)

OpenStudy (lgbasallote):

you took the derivative right?

OpenStudy (ash2326):

3= 2ax+b at (1,1)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

okay so we got 3 -2 = 2a + b

OpenStudy (lgbasallote):

this is substitution of (1,1)?

OpenStudy (anonymous):

1 = 2a +b okay so we take one of them as eqn

OpenStudy (anonymous):

yep

OpenStudy (lgbasallote):

hey...go slow will you -_- lol

OpenStudy (anonymous):

so b = 1- 2a

OpenStudy (lgbasallote):

then..

OpenStudy (ash2326):

@moha_10 Derivative of the parabola at (1,1 ) is the slope of the parabola and tangent to it at (1, 1) We are given \(y=ax^2+bx\) it's slope at (1, 1)=>(m= 2a+b\) From the tangent's equation, slope is 3 so \[3=2a+b\]

OpenStudy (anonymous):

and finally sub to main function

OpenStudy (lgbasallote):

wait..what happened...

OpenStudy (anonymous):

okay may i made a mistake keep going ash

OpenStudy (ash2326):